### v2

@@ 4.3 Ablation Studies
Table 8: Ablation of data augmentation in P2P-point and P2P-voxel, respectively on KITTI dataset.
@@ 4.4 Exploration Studies
Comparison of Convergence Curves. Fig. 10 illustrates the convergence speed of our proposed methods and M2Track [42] on the KITTI Car category. Our curves jitter more sightly compared to M2Track, indicating that the proposed models exhibit faster convergence and greater stability. Furthermore, due to P2P-voxel enables explicit part-to-part fusion between the spatial structures of point clouds in consecutive frames, the Success rate reaches 65.8% in 2 epochs.
Figure 10: Comparison of model convergence speed.
5 Limitation and Discussion
As illustrated in Tab. 2 and 3, tracking methods seem to have nearly reached a state of performance saturation on the KITTI [9] dataset. Nevertheless, on the NuScenes [2] dataset, there is still significant room for developing tracker. Our findings suggest that this phenomenon is largely influenced by the different sparsity in the two datasets. We count the sparsity distribution in the Car category of both the KITTI and NuScenes datasets, as shown in Fig. 11. Moreover, tracking performance corresponding to different sparsity levels is reported in Tab. 10.
@@ 5 Limitation and Discussion
Figure 12: An instance of failure in the proposed PCMT-voxel. We plot input point cloud, feature map on $XY$ plane and output tracking box. The blue box and red box represent our predicted result and ground truth, respectively.
@@
Limitation. The proposed PCMT framework has significantly advanced 3D single object tracking, outperforming previous state-of-the-art methods by a significant margin. Despite this progress, challenges remain evident. Fig. 12 illustrates a tracking failure of our PCMT-voxel method. Tracking drift occurs when target information becomes obscured due to occlusion or extended distance scanning by the LiDAR sensor. Tab. 10 also reflects this phenomenon. This inspires two considerations for future tracking development:

### v3

@@ 4.3 Comparison with State-of-the-art Trackers
Qualitative Comparison. To visually demonstrate the effectiveness of the proposed models P2P-voxel and P2P-point, some visualization comparisons are conducted against the existing motion-centric method M2Track [54] on the Car, Pedestrian, Van and Cyclist categories from the KITTI [10] dataset. As illustrated in Fig. 7, our methods exhibit more accurate object tracking in scenes with distractors or sparse contexts.
Figure 7: Visual results of our P2P-voxel/-point and M2Track on the point cloud sequences of Car, Pedestrian, Van, and Cyclist categories (from top to bottom). The red points and box are the foreground points and ground truth of targets. The blue, yellow and green boxes denote the prediction results by our P2P-voxel, P2P-point and M2Track, respectively.
4.4 Ablation Studies
@@
[…]
Table 9: Ablation of data augmentation in P2P-point and P2P-voxel, respectively on KITTI [10] dataset.
@@ 4.5 Exploration Studies
Figure 11: Comparison of model convergence speed.
Table 11: Comparison of tracking results with or without incorporating box information [54] on KITTI [10] dataset. Not including box information is the default setting.
Tracker # Box Information Car Pedestrian Van Cyclist Mean
P2P-point 68.8/81.7 62.7/89.1 65.4/80.1 74.8/94.8 66.2/85.4
71.0/83.1 64.2/90.7 68.9/83.4 74.3/94.0 68.0/86.7
P2P-voxel 73.6/85.7 69.6/94.0 70.3/83.9 75.5/94.6 71.7/89.4
75.3/86.9 69.7/93.8 72.1/85.0 75.2/94.8 72.7/90.0
@@
Comparison of Convergence Curves. Fig. 11 illustrates the convergence speed of our proposed methods and M2Track [54] on the KITTI Car category. Our curves jitter more rapidly compared to M2Track, indicating that the proposed models exhibit faster convergence and greater stability. Furthermore, due to P2P-voxel enables explicit part-to-part fusion between the spatial structures of point clouds in consecutive frames, the Success rate reaches 65.8% in 2 epochs.
Effectiveness of Box Information. Box size information is a valuable feature for enhancing tracking performance, as demonstrated in BAT [53] and M2Track [54]. Here, we incorporate box information into our proposed P2P-point and P2P-voxel models to evaluate its impact. Following BAT and M2Track, we calculate the pairwise Euclidean distance between each point and the 9 key points of the bounding box. As shown in Tab. 11, incorporating box information significantly improves tracking performance, particularly for the car and van categories. This phenomenon can be attributed to the robustness and accuracy of box information for grid-like objects compared to non-grid objects. Additionally, we observe a larger performance gain in P2P-point than in P2P-voxel. This is because the voxel representation used in P2P-voxel is inherently more effective for handling rigid objects, and its performance has already reached saturation on the KITTI dataset.
@@
Figure 13: An instance of failure in the proposed P2P-voxel. We plot input point cloud, feature map on $XY$ plane and output tracking box. The The blue box and red box represent our predicted result and ground truth, respectively.
5 Limitation and Discussion
As illustrated in Tab. 3 and 4, tracking methods seem to have nearly reached a state of performance saturation on the KITTI [10] dataset. Nevertheless, on the NuScenes [2] dataset, there is still significant room for developing tracker. Our findings suggest that this phenomenon is largely influenced by the different sparsity in the two datasets. We count the sparsity distribution in the Car category of both the KITTI and NuScenes datasets, as shown in Fig. 12. Moreover, tracking performance corresponding to different sparsity levels is reported in Tab. 12.
@@ 5 Limitation and Discussion
Limitation. The proposed P2P framework has significantly advanced 3D single object tracking, outperforming previous state-of-the-art methods by a significant margin. Despite this progress, challenges remain evident. Fig. 13 illustrates a tracking failure of our P2P-voxel method. Tracking drift occurs when target information becomes obscured due to occlusion or extended distance scanning by the LiDAR sensor. Tab. 12 also reflects this phenomenon. This inspires two considerations for future tracking development:
@@ 6 Conclusion
Acknowledgment This work was partially supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT)(RS-2020-II201373, Artificial Intelligence Graduate School Program (Hanyang University)), and the Zhejiang Provincial Key Laboratory of Intelligent Vehicle Electronics Research.